CH/DOC/92/11

Index

Information about the ANU

Author: W.D.M. Rae

Date: 21 Sep 1991

Information About Running at the Australian National University

Information gathered during the September 1992 visit is collected together into a blue A4 book and a folder of loose documents, plus two booklets of photographs.

These are now held at Birmingham, and are indexed below.

Staff

Position  Name
Secretary  Mrs Anne Gillard
Boss  George Dracoulis  academic
Boss  Trevor Ophel  academic
Boss  David Weisser  academic
Targets  Bob Turkentine
Data Acquisition  Gordon Foote
Chamber  Jack Leigh  academic
Seminars  David Hinde  academic
Gas Detectors  Keith Fifield  academic
Machine  David Weisser  academic
Head of Department George Dracoulis  academic
Electronics  John Kennedy  Res. School

Contents of Blue Book

Item Description  Pages
1 Chamber  1—10
2 Workshop  11
3 Minimum Lower Angle  12
4 Cable run  13
5 Control room layout  14
6 Oscilliscope  15
7 Chamber controls  16—18
8 Running in chamber  19
9 Data acquisition  20—24
10 Electricity  25
11 Effort  26
12 Targets  27
13 Internet  28
14 Beam height  29
15 Cabling chamber  30
16 Running dates  31
17 Coffee — tea — DLW  32
18 Isobutane  33
19 Vacuum  34
20 Machine/ beams  35
21 Exchange scheme  36
22 Other documents  37
23 Staff  38
24 University Housing (see also 44)  39
25 Car rental  40
26 Electronics/ cables pool  41
27 Control room facilities  42
28 Drawings  43
29 Housing (see also 39)  44
30 Posted reports  45
31 Sydney Bus  46
32 YHA  48
33 Cars  49
34 E-mail  50
35 Cables, feedthroughs, arm mounts 51—52
36 Notes on final discussions  53

Other Documents

Item Description
1 Annual Report 1991
2 Campus map
3 Accomodation guides, booklets
4 VAX/VMS account, information, telnet numbers
5 Drawings of vacuum chambers
6 Operating vacuum chamber controls
7 Operation of scattering chamber
8 Plan of LINAC
9 Section through scattering chamber
10 Top view of scattering chamber
11 Australian Academy of Science — exchange with UK
12 Bilateral Science and Technology DITCC
13 Cleaning and plating methods for vacuum
14 Vacuum Notes I — III (by post)
15 Mail from Paddy Regan
16 Sample schedule
17 Bill does not believe in number 17
18 Photographs (2m scattering chamber)
19 University Housing Applications
20 Mail from Paddy (2)
21 Bill does not believe in number 21
22 Liversidge Court
23 University House
24 Invitation Letter Draft
25 THK (inner bearings) — addresses
26 Agreement document
27 Isolation of chamber — picture
28 Information on Sydney
29 Map of Canberra
30 Map of Sydney
31 YHA information
32 Original map of campus
33 ANU telephone directory
34 Ursula College
35 Canberra Travellers’ Guide
36 Canberra guide
37 Royal Society Fellowships/ Australia
38 Australian Visa Applications

Drawings

Item Description

2m Scattering Chamber

1 Scattering Chamber Section (1” = 1 cm?)
2 Bolt Holes and Support Fix
3 Target Assembly Mechanism
4 General assembly: Arm Drive
5 Lid matching details
6 Monitor Counter Holder
7 Monitor Counter Plate
8 4–detector angle plate
9 Drive and lead through
10 Details of port locations and sizes
11 Collimator mounting bracket
12 Drive leadthrough
13 Top flange and sight
14 Main drive details
15 Beam layout and Control Room plan

2m Scattering Chamber

16 Monitor Counter Mounts
17 Feedthrough ports on base of chamber

Photgraph Booklet 1 — Scattering Chamber —

No. Description
1 2m Scattering Chamber
2 2m Scattering Chamber — other angle
3 Feedthroughs on base
4 Lid – vacuum interchange
5 View from switcher
6 Local electronics and control
7 Back of chamber
8 Cryopump and control
9 Feedthrough patch
10 Cable path — chamber
11 Cable path — support
12 Cables into duct
13 View of lid
14 Second view of lid
15 Another view of base
16 Lid raising and support
17 Gas feedthroughs
18 Detector Arm
19 Detector arm closeup
20 beam collimator arm
21 Another arm view
22 Arm movement electronics
23 Another arm view
24 Bare arm
25 Centre of chamber
26 Centre closeup

Photgraph Booklet 2 — Charissa —

No. Description
1 Beam collimator arm
2 Beam collimator and tube
3 Target holder
4 Centre (bare)
5 Working in chamber !
6 Chamber beam line
7 Multi-purpose beam line
8 Closer view — Multi-purpose beam line
9 Other side — Multi-purpose beam line
10 Enge chamber
11 Test chamber
12 Control Room — 1
13 Control Room — 2
14 Electronics opposite
15 Electronics to left
16 Two data acquisition stations
17 Space for Charissa
18 Cables out duct
19 ADC’s
20 Data acquisition — 1
21 CAMAC crate — 1
22 CAMAC crate — 2
23 ADC interface
24 ADC’s again

Collated on behalf of Charissa collaboration by William Rae September 1992 and painstakingly typed by WNC